[figure: 9.7 cm]

EFFECTS OF DOMAIN WALL DYNAMICS IN THE OPTICAL

CONDUCTIVITY OF α-NaV_{2}O_{5}

Mario Cuoco^{a,b}, Peter Horsch^{a}, and Frank Mack^{a} aMax-Planck-Institut für Festkörperforschung, Heisenbergstr. 1, D-70569 Stuttgart, Germany b I.N.F.M. di Salerno, Dip. Scienze Fisiche E.R. Caianiello, I-84081 Baronissi Salerno, Italy

Using finite temperature diagonalization techniques it is shown that the quarter-filled t-J-V model on a trellis lattice structure provides a quantitative explanation of the highly anisotropic optical conductivity of the α-NaV_{2}O_{5} compound. The combined effects of the short-range Coulomb interaction and valence fluctuations of V-ions determine the main absorption and the fundamental gap. Inter-ladder hopping is necessary for the explanation of the anomalous in-gap absorption and generation of spectral weight at high energy. The role of valence fluctuations is explained in terms of the domain wall excitations of an anisotropic 2D Ising model in a transverse field close to criticality.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Mario Cuoco

Address for correspondence : Max-Planck-Institut für Festkörperforschung Heisenbergstr. 1,D-70569 Stuttgart (Germany)

Email address : cuoco@audrey.mpi-stuttgart.mpg.de